write equation in slope - intercept form.
Y-1 = 2/3(X+9)
Y-1 = 2/3X+18/3
Y-1+1 = 2/3X+18/3+1
Y = 2/3X+21/3
Is this correct

Answer by Alan3354(69443)

write equation in slope - intercept form.
Y-1 = 2/3(X+9)
Y-1 = 2/3X+18/3
Y-1+1 = 2/3X+18/3+1
Y = 2/3X+21/3
---------------
y-1 = (2/3)x + 6
y = (2/3)x + 7
-------------
Same as your answer, but 21/3 = 7
Also, 2/3x could be either (2/3)x or 2/(3x). Parentheses are free, use some.